# CPathAgent: An Agent-based Foundation Model for Interpretable High-Resolution Pathology Image Analysis Mimicking Pathologists’ Diagnostic Logic

### v1

@@ 5 Conclusion
In this study, we introduced CPathAgent, an agent-based approach for computational pathology that emulates pathologists’ diagnostic reasoning through strategic examination of high-resolution pathology images. By meticulously replicating critical diagnostic processes—such as zooming in to inspect cellular details, zooming out to assess tissue architecture, and systematically shifting focus between regions of interest, CPathAgent achieves both enhanced interpretability and state-of-the-art performance across diverse pathology tasks. Our comprehensive evaluation across three levels of image scale (patch-level understanding, large regional assessment, and whole slide image analysis) demonstrates that this pathology-mimicking approach, which closely mirrors the trained eye of experienced pathologists navigating tissue specimens, effectively bridges the crucial gap between human diagnostic expertise and computational analysis.
@@ A.1 Details of CPathAgent-Instruct
Figure 6: Overview of the VQA-oriented dataset generation process: from multi-scale description to question generation, question filtering, question-oriented navigation path generation, and VQA-related visual reasoning.
Image Sources: CPathAgent derives its WSI report data from HistGen [57], while the corresponding whole slide image data is sourced from The Cancer Genome Atlas (TCGA). TCGA provides a comprehensive collection of pathology whole-slide images contributed by various participating institutions, encompassing a diverse range of distinct tissue and cancer types. This extensive repository ensures diverse coverage of pathological conditions and features essential for dataset and model development.
@@
1) Region Selection Subset Enhancement: For the region selection subset within CPathAgent-Instruct, we face a challenge as this capability was not developed during earlier training stages (as detailed in Section C), and 5,254 WSIs are insufficient for acquiring this skill. Considering that region selection is manageable for advanced large models like Gemini-2.5-Pro and doesn’t heavily depend on WSI reports, we expand our data by utilizing all 24,429 TCGA overview images (excluding test data). The WSI overview images are 32× downsampled versions of the original WSIs as mentioned in the main paper, providing a overview the entire slide at lower resolution. To be specific:
@@
To increase the diversity of our data, we use a generation temperature of 0.8 and create two separate region selection results for each WSI. This strategy produces slightly different region selections from the same WSI overview, effectively doubling our dataset size. At the same time, it introduces natural variations in how regions are selected, which reflects the inherent subjectivity in how different pathologists might evaluate the same slide. The specific prompt templates we use can be found in Section D.5.
2) Navigation Path Diversity Augmentation: Similarly, for the navigation path planning subset, we recognize that pathologists can follow multiple valid paths to reach accurate diagnoses when examining a large region. To capture this diversity, we set the generation temperature to 0.8 and prompt Gemini-2.5-Pro 1-3 times per WSI using the approach described in Section 3.2. This generates diverse navigation paths for the same region. Correspondingly, the multi-scale multi-image reasoning also changes based on the different navigation paths, creating additional variations in our dataset. As a result, we substantially expand our dataset while improving the model’s ability to generate diverse, valid outputs.
3) VQA-oriented Subset Construction and Quality Control: We systematically construct the VQA subset as shown in Figure 6. First, we use 21 multi-scale descriptions from sub-patches at different magnifications (1x, 2x, 4x), which are generated through WSI-report guided prompting (detailed in Section 3.2). We then input these descriptions along with their corresponding region images to Gemini-2.5-Pro to create the initial VQA pairs.
@@
Figure 7: Illustration of the proportional distribution of each subset within the CPathAgent-Instruct dataset.
Figure 8: Illustration of the distribution of regions across tissue types.
Figure 9: Illustration of the distribution of regions across tissue types.
@@
For questions that pass this validation, we generate navigation plans by providing Gemini-2.5-Pro with the question, multi-scale image descriptions, region descriptions, and full region images. Similar to our region description reasoning generation, we combine the validated questions with cropped patches from the planned navigation path and region descriptions. This prompts Gemini-2.5-Pro to generate comprehensive reasoning that systematically addresses each question—progressing logically from overall tissue architecture to specific diagnostic features, comparing and eliminating answer options through continuous reasoning to arrive at the correct conclusion.
@@
Through this question-guided navigation planning and visual reasoning approach, we create a dataset that enables CPathAgent to adapt to diverse diagnostic queries, which allows the model to focus on relevant regions and features based on the specific question.
@@
Proportion of CPathAgent-Instruct subsets Figure 7 illustrates the distribution of different subsets within the CPathAgent-Instruct dataset. Among them, "Viewing Path Planning" and "Multi-Image Understanding and Reasoning" each constitute 32.8% of the total dataset. These equal proportions exist because each planning instance directly corresponds to a reasoning instance, creating a one-to-one relationship between these components. The VQA-oriented subset, though smaller, is sufficient for model adaptation as it shares structural similarities with the main components.
@@
Statistics of Source Regions: As shown in Figure 8, CPathAgent dataset encompasses 26 distinct tissue types with a total of 78,658 individual regions. Breast tissue represents the highest proportion at 13.4% (10,521 samples), while lymphoid tissue appears with 270 samples.
@@
Statistics of Overview WSIs within region selection subset: As shown in Figure 9, the CPathAgent dataset consists of 26 distinct tissue types with a total of 24,429 WSI overviews. Lung tissue represents the highest proportion at 12.9% (3,150 WSIs), followed by Breast with 3,110 WSIs and Kidney with 2,672 WSIs.
@@
Figure 10: Visualization of a word cloud derived from general description-based and VQA-oriented reasoning.
@@
Figure 11: Visualization of a word cloud derived from questions in VQA-oriented subsets in CPathAgent-instruct and PathMMU-HR².
@@
Word Frequency in Multi-scale Multi-image Reasoning Subsets: We analyze word frequency distributions in our General description-based (left, Figure 10) and VQA-oriented (right, Figure 10) reasoning approaches. The visualization reveals distinctive reasoning patterns in pathology diagnosis. General description reasoning emphasizes procedural terms like "zoom," "move," and "examine"—simulating a pathologist’s physical navigation through tissue samples as they methodically adjust magnification, reposition focus areas, and explore specimens. Meanwhile, VQA-oriented reasoning prioritizes diagnostic decision terms like "look," "see," and "feature," reflecting hypothesis testing and evidence evaluation in clinical reasoning. Both approaches incorporate critical pathological terminology ("stroma," "nuclear," "infiltrate") essential for accurate diagnosis, demonstrating that these reasoning strategies center on key diagnostic elements fundamental to clinical pathology assessment and disease classification.
@@
Word Frequency of Questions in VQA-oriented Subset: We analyze the question terminology in our VQA-oriented subset, as shown in the left part of Figure 11. The visualization reveals "diagnosis" as the central focus, surrounded by practical terms like "describe," "integrate," and "feature." Magnification terms ("40x," "10x") indicate how pathologists navigate between different viewing scales, while specific organs ("thyroid," "breast," "lung") appear frequently in questions. Action words like "evaluate" and "observe," along with uncertainty terms like "likely" and "consider," demonstrate the methodical reasoning process. This word pattern demonstrates that our constructed diagnostic questions demand multiscale examination with systematic feature assessment.
@@ A.2 Details of PathMMU-HR²
The construction of PathMMU-HR² is similar to the construction of the VQA-oriented subset in CPathAgen-Instruct, as shown in Figure 6. However, they differ in that after filtering VQA pairs that can be solved through text-only shortcuts, we engage three pathologists with over 10 years of experience for human validation. The evaluation criteria include: clinical relevance of questions, accuracy of provided answers, necessity of multi-scale integration for answering, and alignment with standard pathology practice. Questions failing any criterion are deemed invalid. Under pathologists’ evaluation, we identify 1,688 valid questions from the initially annotated 2,200 questions, with the annotation platform interface shown in Figure 12.
We also visualize the question part of PathMMU-HR² with a word cloud in Figure 11. As shown, the most prominent terms—such as "diagnose," "integrate," "accurately," and "across"—highlight that many queries require multi-scale analysis and complex reasoning. The presence of micro-level descriptors like "nuclear," "cellular," and "morphology," alongside macro-level terms such as "tumor," "lesion," and "pattern" underscores the necessity of integrating observations across different magnification levels (indicated by terms like "40x," "10x," and "2x"). This multi-scale integration is essential for correctly answering questions that demand the model to synthesize information from subcellular details to tissue architecture, evaluating both cytological features and architectural patterns simultaneously to arrive at accurate answers.
Figure 12: Example of the PathMMU-HR2 annotation interface.
@@ B.1 Comparison of Description Generation Performance Among Pathology-specific LMMs
Results: CPathAgent significantly outperforms both PathGen-LLaVA and Quilt-LLaVA across all evaluation metrics. As shown in Figure 13, most notably, CPathAgent achieves a perfect 100% win rate against Quilt-LLaVA across all three evaluators. Against PathGen-LLaVA, CPathAgent maintains consistently strong performance with over 95% win rates across all evaluators.
@@
Figure 13: Comparison of CPathAgent with different pathology-specific LMMs for description/diagnosis generation on 100 randomly sampled cases, evaluated by Gemini-2.5-Pro, GPT-4.1, and expert pathologists.
B.2 Comparison of WSI Classification Performance Among Pathology-specific LMMs
We compare CPathAgent with existing pathology-specific LMMs on WSI classification tasks, following the same approach as used for general LMMs detailed in Section 4.3.
Results: Both PathGen-LLaVA and Quilt-LLaVA significantly underperform on WSI classification, achieving only 60.2% and 54.8% overall balanced accuracy, respectively, as shown in Table 4. This performance gap can be attributed to two key limitations of these models. First, PathGen-LLaVA and Quilt-LLaVA are specifically designed and trained for patch-level analysis using small tissue regions. Consequently, they lack the ability to capture high-level, contextual diagnostic information that is crucial for WSI analysis. Second, these models do not possess the capability to navigate and zoom within large-scale regions. When large regions are resized to fit their input requirements, critical diagnostic details become virtually indiscernible, severely compromising their performance on WSI classification tasks.
Table 4: Overall results of WSI classification tasks. The best-performing model in each subset for general and pathology domain is in-bold, and the model with the second-best performance is underlined. Balanced accuracy (%) is employed.

### v2

@@ 4.4 Out of Distribution Evaluation
While current OOD performance may not yet meet clinical deployment standards, these results demonstrate meaningful generalization beyond TCGA with remarkable data efficiency, suggesting that scaling to larger datasets represents a promising direction for further improvement.
@@ 5 Conclusion
In this study, we present CPathAgent, an agent-based framework for computational pathology that emulates pathologists’ diagnostic reasoning through strategic examination of high-resolution pathology images. Through a systematic three-stage process of global screening, navigation planning, and multi-scale sequence reasoning, CPathAgent faithfully replicates the critical diagnostic workflow by dynamically zooming in and out while systematically shifting focus across regions of interest, transparently documenting not only what features are observed but also why specific regions warrant closer examination and how evidence accumulates across viewing scales, thereby achieving both superior interpretability and state-of-the-art performance across diverse pathology tasks. Our comprehensive evaluation spanning three scales of image analysis (patch-level understanding, large regional assessment, and whole slide image analysis) demonstrates that this agent-based paradigm, which closely mirrors how experienced pathologists navigate and scrutinize tissue specimens, effectively bridges the crucial gap between human diagnostic expertise and AI-assisted analysis.
6 Acknowledgements
This study was partially supported by "Pioneer" and "Leading Goose" R&D Program of Zhejiang (Grant 2025SDXHDX0003), the National Natural Science Foundation of China (Grant No.62506306), foundation of Muyuan Laboratory (Program ID: 14106022401,14106022402), and the Westlake Education Foundation.
@@ A.1 Details of CPathAgent-Instruct
Figure A1: Overview of the VQA-oriented dataset generation process: from multi-scale description to question generation, question filtering, question-oriented navigation path generation, and VQA-related visual reasoning.
Data Sources: CPathAgent derives its WSI report data from HistGen [49], while the corresponding WSI data is sourced from The Cancer Genome Atlas (TCGA). TCGA provides a comprehensive collection of pathology WSIs contributed by various participating institutions, encompassing a diverse range of distinct tissue and cancer types. This extensive repository ensures diverse coverage of pathological conditions and features essential for dataset and model development.
@@
1) Global Screening Subset Enhancement: For the region selection subset within CPathAgent-Instruct, we face a challenge as this capability was not developed during earlier training stages (stage 1 and stage 2 as detailed in Section 3.4), and 5,254 WSIs are insufficient for acquiring this skill. Considering that region selection is manageable for advanced large models like Gemini-2.5-Pro and doesn’t heavily depend on WSI reports, we expand our data by utilizing all 24,429 TCGA overview images (excluding test data). The WSI overview images are 32× downsampled versions of the original WSIs as mentioned in the main paper, providing a overview the entire slide at lower resolution. To be specific:
@@
To increase the diversity of our data, we use a generation temperature of 0.8 and create two separate region selection results for each WSI. This strategy produces slightly different region selections from the same WSI overview, effectively doubling our dataset size. At the same time, it introduces natural variations in how regions are selected, which reflects the inherent subjectivity in how different pathologists might evaluate the same slide. The specific prompt templates we use can be found in Section E.5.
2) Navigation Path Diversity Augmentation: Similarly, for the navigation path planning subset, we recognize that pathologistsmay follow multiple valid trajectories when examining a large region to reach accurate diagnoses. To capture this inherent diversity, we set the generation temperature to 0.8 and prompt Gemini-2.5-Pro randomly 1-3 times per WSI using the approach described in Section 3.2. This stochastic prompting process generates diverse navigation paths for identical regions. Correspondingly, the multi-scale multi-image reasoning also changes based on the different navigation paths, creating additional variations in our dataset. As a result, we substantially expand our dataset while improving the model’s ability to generate diverse, valid trajectories.
3) VQA-oriented Subset Construction and Quality Control: We systematically construct the VQA subset as shown in Figure A1. First, we use 21 multi-scale descriptions from sub-patches at different magnifications (1x, 2x, 4x), which are generated through WSI-report guided prompting (detailed in Section 3.2). We then input these descriptions along with their corresponding region images to Gemini-2.5-Pro to create the initial VQA pairs.
@@
Figure A2: Illustration of the proportional distribution of each subset within the CPathAgent-Instruct dataset.
Figure A3: Illustration of the distribution of regions across tissue types.
Figure A4: Illustration of the distribution of regions across tissue types.
@@
For questions that pass validation, we proceed to generate navigation plans through a two-stage process. First, we provide Gemini-2.5-Pro with the validated question, multi-scale image descriptions, region descriptions, and full region images to generate a navigation path. Second, following the approach established in our region description reasoning generation, we combine the validated question with (1) cropped patches extracted along the planned navigation path and (2) corresponding region descriptions. This combined input prompts Gemini-2.5-Pro to generate comprehensive reasoning that systematically addresses the question—progressing logically from overall tissue architecture to specific diagnostic features, comparing and eliminating answer options through continuous reasoning to arrive at the correct conclusion.
@@
This question-guided navigation and reasoning framework yields a dataset that enables CPathAgent to dynamically adapt to diverse diagnostic queries by focusing on task-relevant regions and features based on the specific question at hand.
@@
Proportion of CPathAgent-Instruct subsets Figure A2 illustrates the distribution of different subsets within the CPathAgent-Instruct dataset. Among them, "Viewing Path Planning" and "Multi-Image Understanding and Reasoning" each constitute 32.8% of the total dataset. These equal proportions exist because each planning instance directly corresponds to a reasoning instance, creating a one-to-one relationship between these components. The VQA-oriented subset, though smaller, is sufficient for model adaptation as it shares structural similarities with the main components.
@@
Statistics of Source Regions: As shown in Figure A3, CPathAgent dataset encompasses 26 distinct tissue types with a total of 78,658 individual regions. Breast tissue represents the highest proportion at 13.4% (10,521 samples), while lymphoid tissue appears with 270 samples.
@@
Statistics of Overview WSIs within region selection subset: As shown in Figure A4, the CPathAgent dataset consists of 26 distinct tissue types with a total of 24,429 WSI overviews. Lung tissue represents the highest proportion at 12.9% (3,150 WSIs), followed by Breast with 3,110 WSIs and Kidney with 2,672 WSIs.
@@
Figure A5: Visualization of a word cloud derived from general description-based and VQA-oriented reasoning.
@@
Figure A6: Visualization of a word cloud derived from questions in VQA-oriented subsets in CPathAgent-instruct and PathMMU-HR².
@@
Word Frequency in Multi-scale Multi-image Reasoning Subsets: We analyze word frequency distributions in our General description-based (left, Figure A5) and VQA-oriented (right, Figure A5) reasoning approaches. The visualization reveals distinctive reasoning patterns in pathology diagnosis. General description reasoning emphasizes procedural terms like "zoom," "move," and "examine"—simulating a pathologist’s physical navigation through tissue samples as they methodically adjust magnification, reposition focus areas, and explore specimens. Meanwhile, VQA-oriented reasoning prioritizes diagnostic decision terms like "look," "see," and "feature," reflecting hypothesis testing and evidence evaluation in clinical reasoning. Both approaches incorporate critical pathological terminology ("stroma," "nuclear," "infiltrate") essential for accurate diagnosis, demonstrating that these reasoning strategies center on key diagnostic elements fundamental to clinical pathology assessment and disease classification.
@@
Word Frequency of Questions in VQA-oriented Subset: We analyze the question terminology in our VQA-oriented subset, as shown in the left part of Figure A6. The visualization reveals "diagnosis" as the central focus, surrounded by practical terms like "describe," "integrate," and "feature." Magnification terms ("40x," "10x") indicate how pathologists navigate between different viewing scales, while specific organs ("thyroid," "breast," "lung") appear frequently in questions. Action words like "evaluate" and "observe," along with uncertainty terms like "likely" and "consider," demonstrate the methodical reasoning process. This word pattern demonstrates that our constructed diagnostic questions demand multiscale examination with systematic feature assessment.
@@ A.2 Details of PathMMU-HR²
The construction of PathMMU-HR² is similar to the construction of the VQA-oriented subset in CPathAgen-Instruct, as shown in Figure A1. However, they differ in that after filtering VQA pairs that can be solved through text-only shortcuts, we engage three pathologists with over 10 years of experience for human validation. The evaluation criteria include: clinical relevance of questions, accuracy of provided answers, necessity of multi-scale integration for answering, and alignment with standard pathology practice. Questions failing any criterion are deemed invalid. Under pathologists’ evaluation, we identify 1,668 valid questions from the initially annotated 2,200 questions, with the annotation platform interface shown in Figure A7.
We also visualize the question part of PathMMU-HR² with a word cloud in Figure A6. As shown, the most prominent terms—such as "diagnose," "integrate," "accurately," and "across"—highlight that many queries require multi-scale analysis and complex reasoning. The presence of micro-level descriptors like "nuclear," "cellular," and "morphology," alongside macro-level terms such as "tumor," "lesion," and "pattern" underscores the necessity of integrating observations across different magnification levels (indicated by terms like "40x," "10x," and "2x"). This multi-scale integration is essential for correctly answering questions that demand the model to synthesize information from subcellular details to tissue architecture, evaluating both cytological features and architectural patterns simultaneously to arrive at accurate answers.
Figure A7: Example of the PathMMU-HR2 annotation interface.
@@ B.1 Comparison of Description Generation Performance Among Pathology-specific LMMs
Results: CPathAgent significantly outperforms both PathGen-LLaVA and Quilt-LLaVA across all evaluation metrics. As shown in Figure A8, most notably, CPathAgent achieves a perfect 100% win rate against Quilt-LLaVA across all three evaluators. Against PathGen-LLaVA, CPathAgent maintains consistently strong performance with over 96% win rates across all evaluators.
@@
Figure A8: Comparison of CPathAgent with different pathology-specific LMMs for description/diagnosis generation on 100 randomly sampled cases, evaluated by Gemini-2.5-Pro, GPT-4.1, and expert pathologists.
B.2 Pathologist evaluation results
We invited a pathology expert to evaluate both diagnostic description generation and VQA tasks across 40 regions, assessing whether the navigation path planning and multi-scale multi-view reasoning align with clinical logic. The model was run 8 times to calculate pass@k metrics.
Results: Pathologist evaluation confirms that both navigation paths and diagnostic reasoning meet clinical expectations. As shown in Figure A9, the model demonstrates strong performance with both tasks. For navigation path evaluation, CPathAgent’s initial satisfaction rates are acceptable but modest, starting at 70% for general description generation and 77.5% for VQA. However, the pass@k metric demonstrates promising improvement when increasing model runs, with both tasks ultimately exceeding 85% satisfaction. For reasoning evaluation, satisfaction rates start higher at 82.5% for general description generation and 87.5% for VQA, and show continued improvement, reaching 90% for general description generation and 92.5% for VQA at higher k values.
Across both tasks, the consistent upward trend in pass@k metrics suggests that the model can reliably generate clinically sound navigation strategies and reasoning when given multiple attempts. Additionally, VQA-based approaches consistently outperform general description tasks across both navigation planning and reasoning. We attribute this to VQA’s focused attention on target features that enables better identification of key elements, whereas general descriptions require additional judgment about content prioritization.
Figure A9: Success rates (pass@k) across 8 generations, evaluated by pathologists across 40 regions. Results show the percentage of pathologist satisfaction with navigation paths and reasoning for General Description Generation (left) and VQA (right) as the number of generations increases.
B.3 Additional Results on PathMMU-Pro
While CPathAgent demonstrates strong performance on PathMMU, achieving results that exceed the reported expert baselines, we acknowledge that the observed gap between our model and expert performance can be attributed to two primary factors inherent to the benchmark design. First, the expert annotations in PathMMU were conducted by pathologists with specialized expertise in one or two specific domains, whereas the dataset encompasses a broad spectrum of tissue types and pathological conditions across multiple subspecialties. Consequently, no individual expert possessed comprehensive knowledge across all domains represented in the evaluation set. Second, as discovered by the PathMMU, large multimodal models may exploit text-based reasoning shortcuts, leveraging strong language priors and contextual cues from question formulations that are not readily accessible to human experts performing traditional diagnostic tasks.
To address the concern regarding text-based shortcuts and provide a more rigorous assessment of our model’s genuine pathological reasoning capabilities, we conducted additional evaluation on PathMMU-Pro [63], a refined version of the benchmark specifically designed with more carefully constructed questions that minimize textual biases.
Results: On PathMMU-Pro, CPathAgent no longer surpasses expert performance but maintains a substantial lead over the previous SOTA CPath-Omni. As shown in Table A4, expert performance now leads with an average accuracy of 69.4%, surpassing CPathAgent’s results. This shift substantiates that the more rigorous question design in PathMMU-Pro effectively mitigates the text-based guessing capabilities that LMMs could previously exploit. Nevertheless, CPathAgent demonstrates consistent advantages over CPath-Omni, achieving 66.3% versus 62.1% on the Test-Tiny subset and 65.3% versus 61.8% on the Test-All subset, demonstrating that CPathAgent possesses stronger multimodal understanding capabilities for pathology.
Table A4: Overall results of models on the PathMMU-Pro. The best-performing model in each subset is in-bold.
Test Overall PubMed SocialPath EduContent
Tiny ALL Tiny ALL Tiny All Tiny All
(430) (4755) (147) (1865) (136) (1081) (147) (1809)
Expert performance 69.4 - 71.2 - 70.1 - 66.9 -
CPath-Omni 62.1 61.8 60.5 60.8 58.8 62.1 66.7 63.0
CPathAgent (Ours) 66.3 65.3 64.6 65.8 65.4 63.6 68.7 66.2
B.4 Ablation Study on Agent Components
To assess the contribution of each agent component, we conduct systematic ablation studies by isolating individual module impacts. We evaluate two critical components. For global screening, we replace the learned region selection module with random selection of 70% of WSI regions. For navigation planning and multi-scale reasoning, we ablate the dynamic navigation and multi-view analysis by directly generating descriptions from entire regions at once, followed by WSI classification based on these single-view descriptions.
Results: Both agent components contribute meaningfully to performance, with navigation and multi-scale reasoning being particularly critical for diagnostic accuracy. As shown in Table A5, removing strategic global screening leads to a 2.5% performance drop (from 82.8% to 80.3%), demonstrating that learned region selection effectively identifies diagnostically informative areas and mimics pathologists’ selective attention during screening. The impact of removing navigation and multi-scale reasoning is substantially more pronounced, with performance degrading by 7.8% (from 82.8% to 75.0%). This significant drop reveals two key limitations of direct WSI region processing: (1) directly processing high-resolution regions (up to $\times16000$ pixels) necessitates aggressive downsampling to fit model input constraints, causing critical diagnostic details to become indiscernible; (2) the absence of pathologist-like reasoning that strategically examines multiple fields of view at varying magnifications fundamentally limits the model’s diagnostic capability. These results confirm that CPathAgent’s effectiveness stems from the integration of all components, with navigation planning and multi-scale reasoning being particularly essential for achieving high diagnostic accuracy on WSI tasks.
Table A5: Results of ablation study on agent components for WSI classification. Balanced accuracy (%) is reported.
Model TCGA-BRCA TCGA-NSCLC TCGA-RCC TCGA-ESCA TCGA-BLCA TCGA-THCA Avg.
Full CPathAgent 88.5 90.8 94.6 97.1 62.7 63.2 82.8
w/o Gloabal Screening 84.7 88.7 94.6 94.1 60.8 58.9 80.3
w/o Navigation & Reasoning 80.0 83.6 88.7 94.1 52.2 51.3 75.0
Table A6: Results of ablation study on applying Gemini-2.5-Pro as an agent for WSI tasks. Balanced accuracy (%) is reported